OpenStudy (anonymous):

square root of 7 - square root of 5 / square root of 7 + square root of 5

OpenStudy (anonymous):

\[\sqrt{7}-\sqrt{5}\div \sqrt{7}+\sqrt{5}\]

OpenStudy (anonymous):

\[(11-2\sqrt{35})/2\]

OpenStudy (anonymous):

not correct

OpenStudy (anonymous):

\[\frac{\sqrt{7}-\sqrt{5}}{\sqrt{7}+\sqrt{5}}\] \[=\frac{\sqrt{7}-\sqrt{5}}{\sqrt{7}+\sqrt{5}}\times\frac{\sqrt{7}-\sqrt{5}}{\sqrt{7}-\sqrt{5}} \]

OpenStudy (anonymous):

\[=\frac{(\sqrt{7}-\sqrt{5})^2}{2}\]

OpenStudy (anonymous):

sorry its 12-2sqrt35 there

OpenStudy (anonymous):

so it is 6-sqrt35

OpenStudy (anonymous):

\[=\frac{12-2\sqrt{35}}{2}=6-\sqrt{35}\]